# Language Repository for Long Video Understanding

### v1

@@ A.1.0.2 Log-likelihood classifier:
A.2 Qualitative examples of repository entries
We present qualitative examples from EgoSchema [51] dataset to better clarify the operations in LangRepo. In Fig. 4, we show the format of repository entries. Here, non-redundant captions from the input get directly written to the repo. In contrast, any redundant captions— grouped based on similarity— get rephrased as concise descriptions (1 per-group). Each repository description may come with additional metadata such as timestamps and #occurrences to avoid the loss of meaningful information due to pruning. In Fig. A.1, we further elaborate on multiple scales within the repository, which are generated by iteratively processing increasingly-longer chunks (created by re-chunk operation). During reading, we can decide to summarize information at various temporal scales to generate output descriptions useful for VQA.

### v2

@@ A.2.0.2 Log-likelihood classifier:
A.3 Qualitative examples of repository entries
We present qualitative examples from EgoSchema (Mangalam et al., 2024) dataset to better clarify the operations in LangRepo. In Fig. 4, we show the format of repository entries. Here, non-redundant captions from the input get directly written to the repo. In contrast, any redundant captions— grouped based on similarity— get rephrased as concise descriptions (1 per-group). Each repository description may come with additional metadata such as timestamps and #occurrences to avoid the loss of meaningful information due to pruning. In Fig. A.1, we further elaborate on multiple scales within the repository, which are generated by iteratively processing increasingly-longer chunks (created by re-chunk operation). During reading, we can decide to summarize information at various temporal scales to generate output descriptions useful for VQA.